A boulder is moved by applying force to two ropes that are tied to it. The forces 120 lb towards the


west and 150 lb towards south . What is the net force applied to the boulder?

1
Expert's answer
2022-06-02T10:26:28-0400

The net force

"F=\sqrt{F_x^2+F_y^2}=\sqrt{120^2+150^2}=192\:\rm lb"